OpenStudy (anonymous):

Can you help me to find the center and the radio of the following circle. (2x^2)+(2y^2)-8x+12y-28=0

OpenStudy (anonymous):

CENTRE IS (2,-3) radius squareroot(42)

OpenStudy (anonymous):

is dis d ans tel

OpenStudy (anonymous):

dont know. Can you teach me how to? Dont care about the answer

OpenStudy (anonymous):

step by step

OpenStudy (anonymous):

k first divide eqn by 2 so that there is no coefficient of x^2 and y^2

OpenStudy (anonymous):

what if I have a 2 in x^2 and a 3 in y^2?

OpenStudy (anonymous):

i dont thnk thats possible bcoz coefficient of x^2 is always equal to coefficient of y^2 in a circle

OpenStudy (anonymous):

nice. So after dividing the equation by 2 I got: x^2+y^2-4x+y-14=0 what else should i do?

OpenStudy (anonymous):

then compar with eqn x^2 +y^2 +2gfx+2hy+c=0 we see -4x=2gfx n 6y=2hy g=-2 n h=3 centre(-g,-h) (2,-3) n radius r^2=g^2 +h^2-c r^2=4+9+14 r^2=27 r=3\[\sqrt{3}\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

checking it out :) till now yes

OpenStudy (anonymous):

Ok in the third line you state that -4x=2gfx and in the next g=-2. how do you get rid of f?

OpenStudy (anonymous):

srry there is no f remove wherever i hav used

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hapy it helpd u

OpenStudy (anonymous):

According to what I understand, we have to leave x^2 and y^2 with coefficient 1. Is that correct?

OpenStudy (anonymous):

Then we have to match the equation with x^2 +y^2 +2gfx+2hy+c=0.

OpenStudy (anonymous):

find the values of g and h and we get the radius.

OpenStudy (anonymous):

sorry the center